Use present value ordinary annuity or amoritization formula to find N PV =7,000 i=0.025, payment-650 what is n?
media%2F5f2%2F5f28b103-9aaf-47a9-9ea2-b2
Answer #1

Calculating the Payment Amount per Period The formula for calculating the payment amount is shown below. 72 A P 1+r)1 where .

Thus,

650 = (7000 * 0.025 * (1+0.025)^n) / ((1+0.025)^n - 1)

Thus,

650 ((1+0.025)^n - 1) = (7000 * 0.025 * (1+0.025)^n)

650(1.025)^n - 650 = 175*(1.025)^n

Thus, 465(1.025)^n = 650

Thus,(1.025)^n = 650/465

solving we get,

n = 12.70